Status Report

NASA Space Station On-Orbit Status 24 Feb 2004

By SpaceRef Editor
February 24, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 24 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

In a final review of the upcoming Orlan EVA, the IMMT (ISS Mission Management Team) determined crew and station ready to proceed with Thursday’s spacewalk. 

On board, EVA preparations continue on schedule.  Today, after wake-up (5:45am EST), workday started with the crew pulling together all the tools and equipment required for the EVA, as specified by an uplinked list.  The gear was bundled on the standard Russian EVA integrated equipment carrier (KPU).   [Kaleri and Foale tied down generic tools, such as hammer, screwdriver, cutters, etc., in the KPU with bungee cords.  Also secured to the KPU bundle, with Aramid tape and short tethers, were carrying bags with two SKK sample containers, two Kromka contamination collection plates, dry towels for Orlan glove decontamination, and the large Matryoshka package, to minimize the size of the EVA bundle for it to pass through the DC-1 docking compartment’s EVA hatch (max. allowable diameter: 1 m).]

The KPU prep activities were video-recorded by FE Alexander Kaleri.  The imagery will be downlinked to the ground later today (5:22pm) via S- and Ku-band for review by specialists.

CDR/SO Michael Foale collected and moved items that needed to be transferred from the U.S. segment (USOS) to the Russian segment (RS) prior to hatch closure tomorrow night (~7:15pm), which will isolate the USOS for the duration of the EVA.  During this lockout period, Foale will “camp out” in the Service Module(SM)’s second sleep compartment.  Reopening of U.S. hatches is nominally planned for Friday (2/27), at ~12:05pm, but may be done earlier on Foale’s preference.   [Transferred items include a laptop with spare hard drive and cables, CD library disks, procedures books, crew provisions, personal health items, a contingency tool kit with tethered scissors to be used outside for cutting the Velcro strap obstructing the WA2 ham radio antenna, etc.]

At 12:30pm, the crew conducted a 20-min. teleconference with ground specialists to review and discuss EVA emergency procedures.   [The off-nominal unmanned ISS configuration during the EVA-9 requires deviations from the standard emergency procedures (in case of smoke/fire and depress), both for the pre-EVA & post-EVA configuration as well as for the period of EVA ops and crew isolation in the DC-1.  In the former config, the RS is open to itself except the Progress (which is ready for ground-commanded undocking if needed) but isolated from the USOS; in the latter config, the RS is isolated from both the DC-1 and the USOS.]

Progress-260/13P was made ready for contingency undocking.   [Alex Kaleri first activated the cargo vehicle, then uninstalled and removed the air ventilation ducts to the SM.  After the SM thrusters were disabled by ground command (U.S. CMGs remaining in momentum-management control), the crew removed the threaded QD (quick disconnect) screw clamps of the SM’s docking and internal transfer mechanism (SSVP), which rigidized the mating surfaces.  Kaleri inspected the clamps for any damage.  Later, the FE will close the hatches in the vestibule tunnel between SM and the cargo vehicle (~2:50pm) and start the standard one-hour leak check of the vestibule with closed hatches.  SM thrusters will be re-enabled at 4:10pm.]

Sasha powered down the SKR-05 hardware of the Russian TEKh-25 Skorpio (“scorpion”) payload in the SM “floor” underneath the TVIS.   [Skorpio’s objective is to monitor environmental radiation parameters with dosimeters inside station compartments at various places and to characterize environmental conditions for conducting scientific and technical experiments.  To operate, Skorpio requires about 6 W of energy.]

Kaleri downloaded data and imagery collected of the BIO-5 Rasteniya-2 (“Plants-2”) experiment to the computer for later downlink to the ground.   [Rasteniya studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.]

The FE also conducted the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO).

Mike Foale tended the PromISS-3 crystal growth experiment in the Microgravity Science Glovebox, today removing tape #11 and installing tape #12, which will capture the final six days of PromISS data.

The crew worked out on TVIS, RED exerciser and VELO cycle with load trainer, and Mike performed the regular transfer of exercise data files from the TVIS treadmill to the medical equipment computer (MEC) for subsequent downlink.

Early this morning, on Daily Orbit 4, TsUP conducted a remote test of the Regul-OS system’s second string.   [Regul is a subsystem of the onboard radio complex (BRK) in the SM, designed for two-way voice communication, digital command/program data, as well as telemetry transmission via RGS (Russian groundsites).  It also has the capability to receive and transmit range, radial velocity, and time-referenced data.  Its proper function is critically important for the EVA, since it is the nominal uplink path for all Russian commands.]

ISS flight attitude, currently still in solar-oriented XPOP (x-axis perpendicular to orbit plane), will be maneuvered to earth-oriented LVLH XVV tomorrow at 2:36pm EST (local vertical local horizontal, x-axis in velocity vector), in support of Thursday’s EVA.

As reported by the crew, some GLA light fixtures (General Luminaire Assembly) need to be replaced in the USOS.  The ground has requested four new GLAs to be manifested on the next Progress, 14P.   [There are two failed GLAs (also known as LHAs, Light Housing Assemblies), one each in the Airlock and Node, and two, reported as “dim”, in the Node.  To assist with cause determination (e.g., are they end-of-life failures?), photos of the units are requested.]

Elektron Troubleshooting:   During the EVA, the Elektron O2 generator will not be running.  Later tonight, on TsUP Go, Kaleri will tear down the equipment set up yesterday for leak-checking the water feed hose (A-R) between the KOV water container (EDV) and the Elektron-VM.  The test, which takes ~24 hrs, has so far shown no pressure change in the pumped-down second EDV water container, indicating that there is no leak from the feed hose or its connection to the KOV EDV.  Moscow plans two more troubleshooting tests on the Elektron, both post-EVA, viz.: an electrical checkout of its cabling, and a nitrogen purge of the internal lines of the Liquid Unit (BZh).  If both are unsuccessful, the BZh will be replaced around 3/6 with a new unit (#7) from spares.

A 20 mmHg repress with air (not O2) from Progress 13P was performed this morning, raising total cabin pressure to ~750 mmHg (~14.5 psi).  There will be no other repress prior to the EVA, but following the spacewalk, with cabin pressure lowered to an estimated 721-723 mmHg, a 13-15 mmHg O2 repress from Progress is required.  If the Elektron still cannot be used at that time, the crew will have to start burning SFOG candles (two per day if the U.S. performs a 15 mmHg repress with 5-9 kg of nitrogen around 3/9, as recommended by TsUP, or four per day without the N2 repress).  A special Commission has been appointed in Moscow to investigate the root cause of the recent failures of the Elektron, which over the last three days has operated on and off for 51 hrs, providing ~8 mmHg (torr) of ppO2 (oxygen partial pressure).

According to Russian reboost/phasing burn planning to meet Soyuz 8S launch and 7S landing constraints, the first maneuver (of 2.3 m/s delta-V) is scheduled for 3/3, followed by a second burn on 3/31 (~2.5 m/s).  The delay of the Shuttle RTF (return to flight) to 2005 has added to the need for a third reboost in April, to meet station altitude requirements.   [8S launch continues to be set for 4/19, 7S landing for 4/29.]

Today’s CEO targets (excluding Asia due to the current sleep cycle and North American and European targets due to winter weather conditions), constrained in XPOP attitude by flight rule to having the Lab science window only available for ~1/4 of each orbit when not facing in flight (“ram”) direction, were Impact craters, Central Libya (pass just NE of the zone: images looking right will provide background material for more detailed views of these recently discovered, hard-to-see features), Inland deltas, SW Sudan (good pass for imaging landscapes immediately west of the Nile and its Sudd swamp [also an inland delta].  Several smaller rivers exit the humid zone of central Africa and enter the Sahara, generating large cones of sediment that are unmapped), Bahamas (good pass along the length of the island chain.  Coral reefs in particular are requested.  Images of the shallow sea floor “platform” are no longer needed), Internal waves, Nicaragua (sunglint opportunity, right of track, for one of the most interesting internal wave zones on the planet in terms of understanding these features), and Internal waves, Tuamotu Is. (internal waves in this remote part of the world ocean of numerous shallow zones are poorly documented.  Looking right towards the glint point as ISS passed over this double island chain).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:50am EST, 2/13).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is Off.  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.8; ppO2 (mmHg) — 159.6; ppCO2 (mmHg) — 3.7;
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.7.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 23.3.
  • Node:  Pressure (mmHg) — 757.88; temperature (deg C) — 23.8 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 750.87; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 760.07; temperature (deg C) — 25.6; shell heater temp (deg C) — 23.7, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 24.7
  • PMA-2:  Shell heater temp (deg C) — 10.4

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position, with 47 deg bias for drag reduction.
  • SM batteries:  All batteries (8) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4070 kg (8972 lb) as of 2/12/04  [SM(755) + FGB(2656) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:21am EST [= epoch]):

  • Mean altitude — 366.2 km
  • Apogee — 370.5km
  • Perigee — 362.0 km
  • Period — 91.87 min.
  • Inclination (to Equator) —  51.628 deg
  • Eccentricity — 0.0006287
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 110 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 30048

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.